OpenStudy (anonymous):

Can someone help me set up the equation thats where i need help. Mutt and Jeff need to paint a fence. Mutt can do the job alone 5 hours faster than Jeff. If together they work for 24 hours and finish only 7/8 of the job, how long would Jeff need to do the job? alone?

OpenStudy (anonymous):

your numbers don't work out evenly, so it makes me think you are typing the problem wrong. we write it or edit your original and i will solve it for you.

OpenStudy (anonymous):

i copied and paste the problem

OpenStudy (anonymous):

@mathandphysic i tried this repeatedly and got a very strange answer. what did you get

OpenStudy (anonymous):

i didn't do it, the numbers are stupid. i will do it give me a second

OpenStudy (anonymous):

i got an equation of \[24\times (\frac{1}{x}+\frac{1}{x-5})=\frac{7}{8}\] but maybe that is wrong

OpenStudy (saifoo.khan):

Sat u r right.

OpenStudy (saifoo.khan):

i just dont wanna solve, it's toooooo long.

OpenStudy (anonymous):

It's a 27 and 3/7 hour job for the 2 of them together, right?

OpenStudy (anonymous):

i don''t know maybe the equation is wrong. lets say jeff can complete the job in 60 hours then jeff can complete it in 55 hours right? so jeff completes 1/60 of the job in an hour and mutt completes 1/55 of the job. combined their rate is \[\frac{1}{60}+\frac{1}{55}\] and if you multiply the rate by 24 you should get 7/8

OpenStudy (anonymous):

just as an example. that is why i thought you should solve \[24(\frac{1}{x}+\frac{1}{x-5})=\frac{7}{8}\] for x

OpenStudy (anonymous):

but when i do that the numbers are so ugly that it must be wrong. so maybe i made a mistake

OpenStudy (anonymous):

unless it is supposed to be mutt can complete the job 5 times faster but if you copied and pasted then that is not right

OpenStudy (anonymous):

i copied and pasted it i went back and look its right

OpenStudy (anonymous):

i cant get it

OpenStudy (anonymous):

what about it can't you get

OpenStudy (anonymous):

the answer as far as i got with that equation is 7x^2-414x+960=0 does it factor because i dont think it does

OpenStudy (anonymous):

you factored it incorrectly.. hold

OpenStudy (anonymous):

24/x + 24/(x-5) = 7/8 common denominator (48x-120)/(x)(x-5) = 7/8 7x^2 - 384x + 925 = 0 you should get x = 2.525084253 and x = 52.3320586 Of course, use the 52.3320586 answer.

OpenStudy (anonymous):

it says that answer is incorrect

OpenStudy (anonymous):

i know, thats why we didn't want to start the problem. your question is wrong.

OpenStudy (anonymous):

i guess its the program... thanks for trying tho, i'll talk to the professor in class

OpenStudy (anonymous):

i have the same question excepts mine says 25 instead of 24.. can someone please help me

OpenStudy (saifoo.khan):

Can you rewrite?

OpenStudy (anonymous):

Mutt and Jeff need to paint a fence. Mutt can do the job alone 5 hours faster than Jeff. If together they work for 25 hours and finish only 7/8 of the job, how long would Jeff need to do the job alone?

OpenStudy (saifoo.khan):

\[25(\frac{1}{x}+\frac{1}{x-5})=\frac{7}{8}\]

OpenStudy (anonymous):

yeah i got that part

OpenStudy (saifoo.khan):

Then you can solve for x

OpenStudy (anonymous):

im having trying finding the common denominator

OpenStudy (saifoo.khan):

LCD = x(x-5)

OpenStudy (anonymous):

thats what i had, & im still getting the wrong answer

OpenStudy (saifoo.khan):

You are correct, mine's solution is also giving wrong answer.

OpenStudy (anonymous):

im getting 101.935 . what did you get?

OpenStudy (saifoo.khan):

57.47 and 2.38

OpenStudy (saifoo.khan):

But there's something wrong in the equation we are making

OpenStudy (saifoo.khan):

Asking for help.. hold on..

OpenStudy (anonymous):

ok